You need to heat 7.59 g of water from 5.2 ºC to 368 K. How much heat is needed? Give answer in Joules and calories.

_____J


_____cal


1
Expert's answer
2021-01-22T06:24:22-0500

5.2°C = (273 + 5.2)k = 278.2k

Change in temperature = 368k -278.2= 89.8k

q (heat energy) = MCp∆T

When Cp of water =1046J

q = 7.59 g × 1046J × 89.8k

= 712, 934.772Jgk-1 = 713kJgk

When Cp of water =250cal

q = 7.59 g × 250 × 89.8k = 170kCalk